### v1

# Energy barriers for small electron polaron hopping in bismuth ferrite from first principles

Abstract
Evidence from first-principles calculations indicates that excess electrons in BiFeO3 form small polarons with energy levels deep inside the electronic band gap. Hence, -type electronic transport could occur by hopping of small electron polarons rather than by band-like transport. Here, by means of first-principles calculations, small electron polaron hopping in BiFeO3 is investigated. Both bulk BiFeO3 and a typical ferroelectric domain wall, the neutral 71° domain wall, are considered. The latter is included to account for experimental observations of electrical conductivity at domain walls in otherwise insulating ferroelectrics. The object of this study is to shed light on the intrinsic electron conduction in rhombohedral BiFeO3 and the effect of pristine neutral ferroelectric domain walls. The computed energy barriers for small electron polaron hopping are of the order of the thermal energy at room temperature, both in bulk and within the neutral 71° domain wall. The domain wall is found to act as a two-dimensional trap for small electron polarons with a trap depth of about two to three times the thermal energy at room temperature. Based on these findings, the intrinsic -type mobility and diffusion constant in BiFeO3 at room temperature are estimated, and experimental conductivity data for BiFeO3 are discussed.
keywords:
polarons, ferroelectrics, domain walls, density-functional theory, electronic transportSabine Körbel*
Institute of Physical Chemistry and Institute of Condensed Matter Theory and Optics, Friedrich Schiller University, Fürstengraben 1, 07743 Jena, Germany
Email Address: skoerbel@uni-muenster.de
1 Introduction
BiFeO3 may be considered as a ferroelectric version of hematite, Fe2O3. BiFeO3 is a perovskite with a rhombohedral crystal structure at room temperature, see Figure 1(a), a ferroelectric polarization of 100 C/cm2 [1], a bandgap of 2.7–3.0 eV [2, 3, 4, 5, 6, 7, 8], and a magnetic structure close to a collinear G-type antiferromagnet.
Like hematite, BiFeO3 might be suitable for water splitting. Other than hematite, BiFeO3 is ferroelectric and can form ferroelectric domain walls, where electronic properties differ from those inside the bulk. In particular, electrical conductivity has been observed at ferroelectric domain walls, whereas the surrounding bulk material is insulating or semiconducting [9, 10, 11, 12, 13, 14]. Like in hematite [15, 16], there is computational evidence that excess electrons in BiFeO3 form small polarons with energy levels deep inside the band gap [17, 18], and that the small electron polarons are trapped by ferroelectric domain walls [19, 20, 21]. Hence, electronic transport in -doped BiFeO3 could occur by hopping of small electron polarons rather than by band-like transport, and the hopping of electron polarons could require less energy than that needed to promote an electron from the polaron state to the conduction band, see Figure 1(b). Domain walls, if present, could affect electronic transport. Here, first-principles electronic structure calculations based on density-functional theory were performed to determine the energy barriers for electron hopping in the bulk and at one of the prevalent ferroelectric domain walls, the neutral 71° domain wall [22, 23]. The 71° domain wall was chosen because it affects the atomic and electronic structure more strongly than the 109° wall [22, 20], and because it has been reported to be conductive in several studies, such as Ref. [13, 24, 25]. The object of this study is to shed light on the intrinsic -type conduction mechanism in BiFeO3 and on the effect of pristine neutral ferroelectric domain walls. The approach taken here allows to determine the intrinsic mechanism of -type conduction, once excess electrons have been brought into BiFeO3 in some way. For example, charge carriers may originate from point defects, such as oxygen vacancies or aliovalent dopants or impurities, from photoexcitation resulting in bound or dissociated electron-hole pairs, or from an electrical contact. The approach taken here can predict the intrinsic charge carrier mobility, but it cannot predict the charge-carrier density since the latter depends on the sample fabrication and history and the experimental conditions, and it can hence not predict the absolute electrical conductivity, but it can predict the temperature dependence of the electrical conductivity and the effect of deviations from the perfect crystal structure. External processes of emitting and trapping electrons by defects are not part of the intrinsic electrical conduction mechanism, and considering the effects of all possible defects would be beyond the scope of this work. However, a particular defect (or desired structure modification) is considered here, namely the neutral ferroelectric 71° domain wall.
This paper is organized as follows: First the computational methods are introduced, then the energy barriers for small polaron hopping are presented for bulk and domain wall and the electronic mobility and diffusion constant at room temperature are determined. Finally results are discussed based on experimental data, where available.
2 Methods
The calculations were performed with the program vasp [26], using the Projector-Augmented Wave (PAW) method [27, 28] and pseudopotentials with 5 (Bi), 16 (Fe), and 6 (O) valence electrons, respectively. Periodic boundary conditions were employed. The local spin-density approximation (LSDA) to density-functional theory (DFT) was used, and the band gap was corrected with a Hubbard- of 5.3 eV applied to the Fe- states using Dudarev’s scheme [29]. This approach yields atomic and electronic structure in close agreement with experiment [19, 20, 21]. Standard approaches to model polarons and polaron hopping from first principles are DFT+ or hybrid DFT functionals, which were applied, e.g., in Ref. [30, 31]. Here, to test how much the calculated hopping barriers depend on the level of theory, the electronic structure with polaron and the polaron hopping barriers in bulk BiFeO3 obtained with LSDA+ and with the hybrid HSE functional [32] for the geometries optimized with LSDA+ are compared. A similar comparison of the hybrid PBE0 functional [33] and DFT+ and a piece-wise linear DFT functional was made in Ref. [34] for polarons in Ga2O3, with the result that polaron formation energy and eigenvalue from DFT+ and PBE0 agree at least semiquantitatively. Plane-wave basis functions with energies up to 520 eV were used. Both the atomic positions and the cell parameters were optimized until the total energy differences between consecutive iteration steps fell below 0.01 meV for the optimization of the electronic density and below 0.1 meV for the optimization of the atomic structure. BiFeO3 with excess electrons was modeled using supercells that comprise eight (bulk) or 16 (domain wall) 10-atom unit cells, respectively. Convergence of energy barriers with respect to supercell size was confirmed for bulk, see Supporting Information. One excess electron was added to the supercell, whose charge was compensated by a uniform background charge density. The magnetic structure of BiFeO3 was approximated by that of a collinear -type antiferromagnet, spin-orbit coupling was neglected.
2.1 Excess electrons in bulk
BiFeO3 bulk with excess electrons was modeled using supercells with 80 atoms that consist of rhombohedral 10-atom unit cells. The first Brillouin zone of the supercell was sampled with -points, equivalent to -points for the primitive rhombohedral ten-atom unit cell of BiFeO3 and to -points for a pseudocubic five-atom perovskite cell with about four Å edge length, as depicted in Figure 2(a).
2.2 Excess electrons at the ferroelectric 71° domain wall
Excess electrons at 71° domain walls were modeled using a 160-atom supercell, see Figure 2(b), spanned by 4() ( in the pseudocubic system, here “”), (- ( in the pseudocubic system), and () ( in the pseudocubic system), where , , and span the pseudocubic 5-atom cell. This supercell is based on that used in Ref. [19, 20, 21]. The atomic structures and formation energies of low-energy ferroelectric domain walls in BiFeO3 are well known [35, 22, 36, 23, 37]. Atomic and electronic structure of the neutral 71° domain wall investigated here have been published elsewhere [23, 22, 19, 20, 21]. The antiferromagnetic -type spin configuration of the bulk was maintained in the systems with domain walls. The first Brillouin zone of the supercell was sampled with -points, equivalent to -points for the primitive rhombohedral ten-atom unit cell of BiFeO3 and to -points for a pseudocubic five-atom perovskite cell with about four Å edge length.
2.3 Polaron hopping
According to previous works [17, 18, 19, 20, 21], the excess electron localizes mostly on a single Fe site. In bulk BiFeO3, all Fe sites are equivalent. Localization on a specific Fe site was induced by means of the occupation matrix control (OMC) method of Allen and Watson [38].
Energy barriers for polaron hopping between different Fe sites were obtained using the nudged-elastic-band (NEB) method, in which an artificial spring force acting along the hopping path constrains atomic structure optimization to directions perpendicular to the path. Nine configurations (NEB images) along each hopping path were considered, including initial and final state.
Both hops between nearest-neighbor (NN) and second-nearest neighbor (2NN) Fe sites were considered, since NN hops require spin flip of the electron and should hence be suppressed.
The number of inequivalent NN and 2NN hops is limited by symmetry. In bulk, there is one NN hop and two inequivalent 2NN hops: a 2NN hop perpendicular to the ferroelectric polarization direction, and a 2NN hop in a direction that forms an angle of ° with the polarization direction, see Figure 2(a). At ferroelectric domain walls, Fe sites are still equivalent within planes parallel to the wall, but differ between these planes. Hops can occur along the wall or with a component perpendicular to the wall, see Figure 2(b). NEB calculations were considered converged when the energy barrier changed by maximally 1 meV in two consecutive iterations of the NEB structure optimization. The strain dependence of the barriers was also investigated. The strain reference (zero strain, ) corresponds to the optimized geometry of bulk or domain wall without excess electron. Tensile strain, , corresponds to the optimized geometry of bulk or domain wall with excess electron.
2.4 Transition state theory
The abrupt jump of the small electron polaron from initial to final site (see below) as a function of the hopping coordinate indicates that the hopping process is non-adiabatic rather than adiabatic. Marcus theory [39] was applied to obtain non-adiabatic hopping rates of small electron polarons between 2NN lattice sites in BiFeO3.
NN hopping, being spin-forbidden, was neglected. The transition rate is then
is the activation energy (hopping barrier) for the 2NN hop, is the reorganization energy. For symmetric hops, . The electronic coupling matrix elements , where the are electronic wave functions localized on the initial and final Fe site, respectively, were adopted from Ref. [40], where they were calculated for Fe(H2O) complexes as a function of the Fe-Fe distance from a model Hamiltonian that treated Fe with a pseudopotential and H2O as classical point charges. Here, the matrix element for the second-nearest Fe-Fe distance calculated for bulk BiFeO3, 5.516 , was used throughout (-101.875 cm eV). Alternatively, the coupling matrix elements could be determined as the deviation of the calculated barrier from the parabolic crossing point, see Figure 3(a). In the case of compressive strain, where the calculated hopping energy profile is close to a double parabola, one obtains a coupling matrix element of to meV.
The electron mobility and the diffusion constant are obtained from the transition rate using the Einstein relation as
where is the elementary charge, is the hopping distance, is the number of equivalent available final states for a given initial state, is the dimension, is Boltzmann’s constant, and is the temperature, here 300 K. For bulk, =3, =12 for 2NN hops, is the 2NN Fe-Fe distance, and the pseudocubic lattice constant (about 4 Å). For two-dimensional transport within a domain-wall plane, . Hops along the coordinate consist of two consecutive hops, such as , with and . Hops along the coordinate can be single 2NN hops with and =2, such as , or double 2NN hops with and , such as . The effective barrier for double hops was taken as the higher one of the two. The two-dimensional -type mobility and the diffusion constant in the domain wall are computed as
3 Results
3.1 Excess electrons in bulk
Electronic structure from LSDA+ and HSE
The electronic density of states (DOS) for a supercell (80 atoms) with an electron polaron is shown in Figure 3(b). The DOS obtained with LSDA+ and HSE agree semiquantitatively.
The computed polaron level lies 0.7 eV below the conduction band minimum for LSDA+ and 1 eV below the conduction band minimum for HSE,
with an estimated uncertainty of 0.2 eV.
[…]
The energy barriers for NN hops and the two different 2NN hops are depicted in Figure 4(a) and (b) for different strains . Both NN and 2NN barriers are of the order of the thermal energy at room temperature and shrink upon compression. Hence, the hopping barrier is an order of magnitude smaller than the energy distance to the conduction band minimum. The 2NN barriers are not strongly direction dependent. The transitions from initial to final site are non-adiabatic (abrupt), as seen from the isosurface of the excess electron’s density as a function of the hopping coordinate (yellow orbitals in Figure 4(c)–(f)). The computed energy barriers are listed in Table 1.
NN | 2NN, ° | 2NN, ° | ||||
LSDA+ | HSE | LSDA+ | HSE | LSDA+ | HSE | |
% | 8 | 6 | 17 | 18 | 16 | 14 |
0 | 15 | 18 | 24 | 34 | 20 | 20 |
0.64% | 15 | 25 | 27 | 39 | 29 | 30 |
To obtain the electron mobility and the diffusion constant at room temperature, the barrier at 0.65%, corresponding to the experimental lattice constant at room temperature [41], was used. With an energy barrier of 28 meV (the average of the two 2NN barriers) and a hopping distance of the 2NN Fe-Fe distance, the computed electron mobility in BiFeO3 bulk at room temperature is 0.6 cm2/Vs, the computed diffusion constant is 0.02 cm2/s. The uncertainty of the computed and are about one order of magnitude.
3.2 Excess electrons at domain walls
When starting from a delocalized-electron configuration, a small electron polaron forms spontaneously in the plane labeled “1” (on the right-hand side of the wall, if the net polarization points to the right, see Figure 2(b)), which is a metastable site. The most stable site is inside the domain wall, in the plane labeled “0”.
Figure 5 shows the energy barriers for electron hops at the 71° domain wall. Only the lowest barriers in each direction are shown. All barriers are depicted and tabulated in the Supplemental Information. The hops are projected on the symmetry-adapted coordinates (see Figure 5) [ wall, (a)], [within the wall, (b)], and [within the wall, , (c)]. For hops between the two trap states in the planes “0” and “1”, the barrier [bright symbols and lines in (a)] is similar to the thermal energy at room temperature, whereas for hops out of the wall (from plane “0” or “1” to plane “-1”,“2”, or “-2”) the barrier is about 2–3 times larger. For hops within the domain wall [(b) and (c)] there are pathes available with barriers similar to the thermal energy at room temperature. Barriers are invariant with respect to translation by a lattice vector parallel to the wall. Like in bulk, the barriers shrink upon compression. The lowest barriers for NN and 2NN hops are similar in size. The computed 2NN energy barriers are listed in Table 2 and in the Supporting Information.
double 2NN, | 2NN, | 2NN, escape | |
or | from DW | ||
% | 31 | 16 | 57 |
0 | 32 | 25 | 62 |
0.34% | 33 | 27 | 65 |
0.65% (extrapol.) | 35 | 29 | 67 |
The computed two-dimensional electron mobility inside the domain-wall plane is cm2/Vs, the computed two-dimensional diffusion constant inside the domain wall is cm2/s. and have an uncertainty of about one order of magnitude.
4 Discussion
Hall measurements for hematite resulted in electron mobilities of about 0.01 to 1 cm2/Vs [42, 43, 44], near the upper end of which lies the here computed mobility in bulk BiFeO3 of 0.6 cm2/Vs. Given the similarities in atomic and electronic structure of BiFeO3 and hematite, it seems plausible that intrinsic mobilities of BiFeO3 and hematite should be comparable.
Experimentally determined activiation energies of electron conduction in BiFeO3 are typically much larger (between about 0.1 and 1.3 eV [45, 46, 47, 48, 49]) than the here computed barrier of about 28 meV, indicating that the experimentally determined barriers do not originate from electron polaron hopping in pristine bulk nor from escape of electron polarons from pristine neutral ferroelectric domain walls, assuming that the other typical walls, the 109° and the 180° walls, exhibit a similar or smaller trap depth than the 71° wall. It seems likely that the experimentally determined barriers of several hundred meV originate from defects other than pristine neutral ferroelectric domain walls, for example from point defects, such as oxygen or bismuth vacancies, or from interface effects, such as Schottky barriers. In Refs. [13, 50] an activation barrier of eV was found for -type conduction, which is about the energy difference between the electron polaron level and the conduction band minimum computed here. In principle, this activation barrier of 0.7 eV could originate from emission of electrons from the polaron state to the conduction band. This would mean that transport by emission from polarons into delocalized conduction states and subsequent ballistic transport would be more efficient than polaron hopping, inspite of the much higher energy barrier that must initially be overcome. However, there are reasons to believe that the lifetime of thermalized electrons in the conduction band at room temperature should be small. This is because the strong Fe- contribution to the states at the conduction band minimum and the small size of the electron polaron indicate that electron-phonon coupling should be large, which should lead to strong electron-phonon scattering. First-principles molecular dynamics simulations yield a polaron formation time of less than a picosecond [20]. Once a small electron polaron has formed, it is unlikely to transform into a delocalized conduction-band electron at room temperature since the energy distance between the polaron level and the conduction-band minimum of eV is much larger than the polaron hopping barrier of meV. It seems most likely that at least for thermalized electrons at temperatures room temperature, small electron hopping should be the dominant intrinsic -type transport process, and that in the reported experimental measurements of the electrical conductivity, the small barrier of meV for electron polaron hopping is hidden behind much larger barriers associated with emission from defects or interfaces other than pristine neutral ferroelectric domain walls.
5 Summary and conclusion
Energy barriers for small electron polaron hopping in BiFeO3 were determined using density-functional theory. Hopping in perfect, pristine BiFeO3 and at pristine, neutral ferroelectric 71° domain walls was considered. In both bulk and domain-wall plane, the calculated energy barriers are of the order of at room temperature ( meV) and are reduced by compressive strain. Using non-adiabatic Marcus theory, an electron mobility of cm2/Vs and cm2/Vs for bulk and within the domain-wall plane, respectively, is obtained at room temperature. The computed diffusion constant in bulk at room temperature is cm2/s both in the bulk and in the domain-wall planes. and have an uncertainty of about one order of magnitude. The domain wall planes act as traps for small electron polarons, with a trap depth of about two to three times at room temperature ( meV). These findings indicate that (a) the effect of pristine, neutral ferroelectric domain walls on the -type conductivity at room temperature consists in accumulating excess electrons and confining their motion to the domain-wall plane. In the Drude picture (conductivitycarrier densitymobility), the electron density along the wall is enhanced compared to the bulk, the mobility perpendicular to the wall is reduced compared to bulk. This is in line with experimental observations that conductivity is enhanced at ferroelectric domain walls and in directions parallel to the domain walls, (b) experimentally found activation barriers of several hundred meV for electrical conductivity in BiFeO3 should have an origin other than electron polaron hopping in pristine BiFeO3 or electron polaron escape from pristine neutral ferroelectric 71° domain walls.
Supporting Information
Supporting Information is available from the Wiley Online Library or from the author.
Acknowledgements
S. K. thanks Michele Reticcioli for recommending the OMC method. This project has received funding from a postdoc stipend of the University of Jena, Germany and from the Volkswagen Stiftung (Momentum) through the project “dandelion”. Computational resources and support were supplied by the HPC cluster ARA of the University of Jena, Germany. Figures were made using Vesta [51] and gnuplot.