### v2

# Energy profile and hopping barriers for small electron polarons at ferroelectric domain walls in bismuth ferrite from first principles

Abstract
Evidence from first-principles calculations indicates that excess electrons in BiFeO3 form small polarons with energy levels deep inside the electronic band gap. Hence, -type electronic transport could occur by hopping of small electron polarons rather than by band-like transport. Here, by means of first-principles calculations, small electron polaron hopping in BiFeO3 is investigated. Both bulk BiFeO3 and a typical ferroelectric domain wall, the neutral 71° domain wall, are considered. The latter is included to account for experimental observations of electrical conductivity at domain walls in otherwise insulating ferroelectrics. The object of this study is to shed light on the intrinsic electron conduction in rhombohedral BiFeO3 and the effect of pristine neutral ferroelectric domain walls. The computed energy barriers for small electron polaron hopping are near 0.2 eV, similar to other perovskite oxides, both in the bulk and within the neutral 71° domain wall. Trapping energies of small electron polarons at the three prevalent domain walls, the 71°, the 109°, and the 180° wall, were determined. The domain walls are found to act as two-dimensional traps for small electron polarons, with a trap depth of about two times the thermal energy at room temperature. Based on these findings, the intrinsic -type mobility and the diffusion constant in BiFeO3 at room temperature are estimated, and experimental conductivity data for BiFeO3 are discussed.
keywords:
polarons, ferroelectrics, domain walls, density-functional theory, electronic transportSabine Körbel*
Institute of Physical Chemistry and Institute of Condensed Matter Theory and Optics, Friedrich Schiller University, Fürstengraben 1, 07743 Jena, Germany
Email Address: skoerbel@uni-muenster.de
1 Introduction
BiFeO3 may be considered as a ferroelectric version of hematite, Fe2O3. BiFeO3 is a perovskite with a rhombohedral crystal structure at room temperature, see Figure 1(a), a ferroelectric polarization of 100 C/cm2 [1], a bandgap of 2.7–3.0 eV [2, 3, 4, 5, 6, 7, 8], and a magnetic structure close to a collinear G-type antiferromagnet.
Like hematite, BiFeO3 might be suitable for water splitting. Other than hematite, BiFeO3 can form ferroelectric domain walls, where electronic properties differ from those inside the bulk. In particular, electrical conductivity has been observed at ferroelectric domain walls in BiFeO3 and in other ferroelectric perovskite oxides, whereas the surrounding bulk material is insulating or semiconducting [9, 10, 11, 12, 13, 14]. Like in the case of hematite [15, 16], there is computational evidence that excess electrons in BiFeO3 form small polarons with energy levels deep inside the band gap [17, 18], and that the small electron polarons are trapped by ferroelectric domain walls [19, 20, 21]. Hence, electronic transport in -doped BiFeO3 should occur by hopping of small electron polarons rather than by band-like transport, if hopping of electron polarons requires less energy than promoting an electron from the polaron state to the conduction band, see Figure 1(b). Ferroelectric domain walls, if present, could affect electronic transport.
Here, for the first time, first-principles electronic structure calculations based on density-functional theory were performed to determine the energy barriers for electron polaron hopping in BiFeO3. Different hopping directions were considered so that the anisotropy of the crystal structure was fully taken into account. For the first time, polaron hopping at a ferroelectric domain wall was modeled from first principles. The neutral 71° domain wall was chosen because it is one of the prevalent ferroelectric domain walls in BiFeO3 [22, 23], because it affects the atomic and electronic structure more strongly than the 109° wall [20, 22], and because it has been reported to be conductive (about one order of magnitude more than the domain interior) in several studies, such as Ref. [13, 24, 25]. The object of this study is to shed light on the intrinsic -type conduction mechanism in BiFeO3 and on the influence of pristine neutral ferroelectric domain walls.
This paper is organized as follows: First the computational methods are introduced, then the energy landscapes and hopping barriers for small polarons are presented for bulk and domain wall, and the electronic mobility and the diffusion constant at room temperature are determined. Finally results are discussed based on experimental data, where available.
2 Methods
The calculations were performed with the “Vienna Ab initio Simulation Package” (Vasp) [26], using the Projector-Augmented Wave (PAW) method [27, 28] and pseudopotentials with 5 (Bi), 16 (Fe), and 6 (O) valence electrons, respectively. Periodic boundary conditions were employed. The local spin-density approximation (LSDA) to density-functional theory (DFT) was used, and the band gap was corrected with a Hubbard- of 5.3 eV applied to the Fe- states using Dudarev’s scheme [29]. This approach yields atomic and electronic structure in close agreement with experiment [19, 20, 21], see also Table 1.
Standard approaches to model polarons and polaron hopping from first principles are DFT+ or hybrid DFT functionals, which were applied, e.g., in Ref. [30, 31]. Here, to test how much the calculated hopping barriers depend on the level of theory, the electronic structure with polaron and the polaron hopping barriers in bulk BiFeO3 from LSDA+ are compared to those from the hybrid HSE06 functional [32] (25% Hartree-Fock exchange, 5 Å screening length) for the geometries optimized with LSDA+. A similar comparison of the hybrid PBE0 functional [33] and DFT+ and a piece-wise linear DFT functional was made in Ref. [34] for polarons in Ga2O3, with the result that polaron formation energy and eigenvalue from DFT+ and PBE0 agree at least semiquantitatively.
Plane-wave basis functions with energies up to 520 eV were used. Both the atomic positions and the cell parameters were optimized until the total energy differences between consecutive iteration steps fell below 0.01 meV for the optimization of the electronic density and below 0.1 meV for the optimization of the atomic structure.
2.1 Excess electrons in bulk
BiFeO3 bulk with excess electrons was modeled using supercells with 80 atoms that consist of rhombohedral 10-atom unit cells. Convergence of energy barriers with respect to the supercell size was confirmed by comparing with results obtained with a 40-atom supercell, see Supp. Inf. The barriers obtained with the two supercell sizes differ only by about 10%. The first Brillouin zone of the supercell was sampled with -points, equivalent to -points for the primitive rhombohedral ten-atom unit cell of BiFeO3 and to -points for a pseudocubic five-atom perovskite cell with about four Å edge length, as depicted in Figure 2(a). Convergence with respect to -point density was confirmed by comparing with results obtained with a mesh, see Supp. Inf. One excess electron was added to the supercell, whose charge was compensated by a uniform background charge density. The magnetic structure of BiFeO3 was approximated by that of a collinear -type antiferromagnet, spin-orbit coupling was neglected.
2.2 Excess electrons at the ferroelectric 71°, 109°, and 180° domain wall
The atomic structures and formation energies (here 172 mJ/m2 for the 71° wall, 63 mJ/m2 for the 109° wall, and 86 mJ/m2 for the 180° wall [20]) of low-energy ferroelectric domain walls in BiFeO3 are well known [22, 23, 35, 36, 37]. Atomic and electronic structure of the neutral 71°, 109°, and 180° domain walls investigated here have been published elsewhere [19, 20, 21, 22, 23]. The computed properties of the ferroelectric 71°, 109°, and 180° domain walls, such as domain-wall energies and widths, can be found in the Supplemental Material to [20]. The antiferromagnetic -type spin configuration of the bulk was maintained in the systems with domain walls. The energy profile of small electron polarons as a function of the distance from the domain walls was determined and the trapping energy at the domain walls was obtained. 240-atom supercells were employed with 12 perovskite layers in the direction perpendicular to the walls (six layers per domain, see Figure 6). -point meshes of -points were used. Very good convergence of results was obtained with this mesh, compared to a -point mesh, see Supporting Information. The energy convergence thresholds were set to eV for the electronic structure and to eV for the atomic coordinates, respectively, which is probably one order of magnitude tighter than necessary. The trapping energy was determined as the total energy difference between a polaron position in the domain interior (DI) and in the lowest-energy configuration at the domain wall (DW), . Energy profiles and trapping energies were corrected by a factor , see Table 1, to avoid underestimation of energy differences that originate in ferroelectricity. Based on the computed trapping energies, the concentration of polarons was estimated based on a Boltzmann distribution,
where is the concentration far away from the domain wall, is the total energy difference between a polaron in and in the domain interior, is Boltzmann’s constant, and is the temperature.
Hopping of small electron polarons at 71° domain walls was modeled using a 160-atom supercell, see Figure 2(b), spanned by 4() ( in the pseudocubic system, here “”), (- ( in the pseudocubic system), and () ( in the pseudocubic system), where , , and span the pseudocubic 5-atom cell. This supercell is based on that used in Ref. [19, 20, 21]. The first Brillouin zone of the supercell was sampled with -points, equivalent to -points for the primitive rhombohedral ten-atom unit cell of BiFeO3 and to -points for a pseudocubic five-atom perovskite cell with about four Å edge length.
2.3 Polaron hopping
According to previous works [17, 18, 19, 20, 21], the excess electron localizes mostly on a single Fe site. In bulk BiFeO3, all Fe sites are equivalent. Localization on a specific Fe site was induced by means of the occupation matrix control (OMC) method of Allen and Watson [38].
Energy barriers for polaron hopping between different Fe sites were obtained using the nudged-elastic-band (NEB) method, in which an artificial spring force acting along the hopping path constrains atomic structure optimization to directions perpendicular to the hopping path. Nine configurations (NEB images) along each hopping path were considered, including initial and final state. The NEB method is designed to determine the minimum-energy path, which can have a considerably lower energy barrier than the shortest path (the linear path). Since here the top of the barrier is pointed (the forces along the minimum energy path are not continuous), which might pose numerical difficulties for the NEB algorithm, the top of the barrier was obtained by calculating the energy along a linear path between the two nearest NEB images (images 4 and 5). The deviation of the barrier from the highest NEB energies obtained for the 2NN hops in bulk were then used to estimate the barrier tops for hops at the domain wall, i.e. the energy of the linear path (between images 4 and 5) was not calculated for the domain-wall hops. Both hops between nearest-neighbor (NN) and second-nearest neighbor (2NN) Fe sites were considered, but NN hops require spin flip of the electron and should hence be suppressed.
The number of inequivalent NN and 2NN hops is limited by symmetry. In bulk, there are one NN hop and two inequivalent 2NN hops: a 2NN hop perpendicular to the ferroelectric polarization direction, and a 2NN hop in a direction that forms an angle of ° with the polarization direction, see Figure 2(a). At ferroelectric domain walls, Fe sites are still equivalent within planes parallel to the wall, but differ between these planes. Hops can occur along the wall or with a component perpendicular to the wall, see Figure 2(b). The NEB calculations were considered converged when the energy barrier changed by maximally 1 meV in two consecutive iterations of the NEB structure optimization. The strain dependence of the barriers was also investigated. The strain reference (zero strain, ) corresponds to the optimized geometry of bulk or domain wall without excess electron. Tensile strain, , corresponds to the optimized geometry of bulk or domain wall with an excess electron.
2.4 Transition state theory
The abrupt jump of the small electron polaron from initial to final site (see below) as a function of the hopping coordinate indicates that the hopping process is non-adiabatic rather than adiabatic. Marcus theory [39] was applied to obtain non-adiabatic hopping rates of small electron polarons between 2NN lattice sites in BiFeO3.
NN hopping, being spin-forbidden, was neglected. The transition rate is then
is the activation energy (hopping barrier) for the 2NN hop, is the reorganization energy. For symmetric hops, . The electronic coupling matrix elements , where the are electronic wave functions localized on the initial and final Fe site, respectively, were adopted from Ref. [40], where they were calculated for Fe(H2O) complexes as a function of the Fe-Fe distance from a model Hamiltonian that treated Fe with a pseudopotential and H2O as classical point charges. Rosso and coworkers obtain a coupling matrix element of 60 meV for second-nearest neighbors in hematite [41], where the intersite distance is smaller and electronic intersite coupling should be larger than in BiFeO3. Here, the matrix element for the second-nearest Fe-Fe distance calculated for bulk BiFeO3, 5.516 , was used throughout (-101.875 cm eV). Alternatively, the coupling matrix elements could in principle be determined as the deviation of the calculated barrier from the parabolic crossing point, see Figure 1(b). In this case one obtains a coupling matrix element of the order of meV, see Supp. Inf., with a large error bar. This result is not used here.
The electron mobility and the diffusion constant are obtained from the transition rate using the Einstein relation as
where is the elementary charge, is the hopping distance, is the number of equivalent available final states for a given initial state, is the dimension, is Boltzmann’s constant, and is the temperature, here 300 K. For bulk, =3, =12 for 2NN hops, is the 2NN Fe-Fe distance, and the pseudocubic lattice constant (about 4 Å). For two-dimensional transport within a domain-wall plane, . Hops along the coordinate consist of two consecutive hops, such as or , with and . Hops along the coordinate can be single 2NN hops with and =2, such as , or double 2NN hops with and , such as or . The two-dimensional -type mobility and the diffusion constant in the domain wall are computed as
3 Results
3.1 Material properties of BiFeO3 obtained with different exchange-correlation functionals
Table 1 contains properties of BiFeO3 in the cubic and the rhombohedral phase obtained with different exchange-correlation functionals and from experiment. The calculations were performed using the primitive rhombohedral 10-atom unit cell of BiFeO3. A cutoff energy of 520 eV, a -point mesh of points, and energy convergence thresholds of eV and eV for electronic and atomic structure were employed, respectively, except for HSE06, where a -point mesh of points was used. The ionic polarization was computed as a weighted sum of atomic displacements,
Here, denotes atoms in the unit cell, and are actual positions in the ferroelectric (FE) and reference positions in the paraelectric (PE) phase, the charges are nominal ionic charges (Bi3+, Fe3+, O2-), the are weighting factors, and is the unit-cell volume.
| (eV) | (Å) | (Å) | (°) | (C/cm2) | |
| HSE06 | -1.15 | 5.48 | 5.63 | 59.15 | 73 |
| LSDA | -0.55 | 5.26 | 5.43 | 60.53 | 63 |
| LSDA+ | -0.90 | 5.44 | 5.52 | 59.82 | 61 |
| PBE | -0.73 | 5.54 | 5.59 | 59.44 | 52 |
| PBE+ | -0.84 | 5.56 | 5.71 | 59.05 | 68 |
| PBEsol | -1.00 | 5.46 | 5.54 | 59.93 | 66 |
| expt. | 5.64 | 59.42 | 72 |
Taking PBEsol or HSE06, which yield excellent structural agreement with experiment, as a benchmark, LSDA+ underestimates the ferroelectric energy gain by about 10–20%.
3.2 Excess electrons in bulk
Electronic structure from LSDA+ and HSE
| LSDA+ | HSE06 | |
| (eV) | -0.34 | -0.57 |
| (eV) | -0.71 | -1.2 |
The formation energy of small polarons is strongly negative (LSDA: 0.3 eV, HSE: 0.6 eV, see Table 2), hence the small polaron is the ground state and should be stable not only at low temperature, but also at room temperature and above. The electronic density of states (DOS) for a supercell (80 atoms) with an electron polaron is shown in Figure 3. The DOS obtained with LSDA+ and HSE06 agree semiquantitatively. The computed polaron level lies 0.7 eV below the conduction band minimum for LSDA+ and 1 eV below the conduction band minimum for HSE06, with an estimated uncertainty of 0.2 eV. Hence the polaron level is a deep level at room temperature.
The moderate strain dependence of polaron formation energy and electronic eigenvalue is depicted in Figure 4.
Hopping barriers
The energy barriers for NN hops and the two different 2NN hops are depicted in Figure 5(a) and (b) for different strains . Both NN and 2NN barriers are about 0.2 eV and slightly increase under tensile strain. Hence, the hopping barrier is considerably smaller than the energy distance of about 0.7 eV to the conduction band minimum. The 2NN barriers are not strongly direction dependent. The transitions from initial to final site are non-adiabatic (abrupt), as seen from the isosurface of the excess electron’s density as a function of the hopping coordinate (yellow orbitals in Figure 5(c) and (d)). The computed energy barriers are listed in Table 3.
| NN | 2NN, ° | 2NN, ° | ||||||
| LSDA+ | HSE06 | LSDA+ | HSE06 | LSDA+ | HSE06 | (cm2/Vs) | (cm2/s) | |
| % | 0.15 | 0.23 | 0.16 | 0.25 | 0.16 | 0.24 | – | – |
| 0 | 0.16 | 0.19 | 0.18 | 0.20 | 0.17 | 0.21 | – | – |
| 0.64% | 0.17 | 0.26 | 0.18 | 0.22 | 0.18 | 0.22 | ||
| 1.9% | 0.19 | – | 0.18 | – | 0.20 | – | ||
The barrier obtained with HSE06 is considerably larger, however the structure was not optimized with HSE06 but with LSDA+, therefore the barrier from HSE06 should be overestimated. In the following, the barrier obtained with LSDA+ is used. To obtain the electron mobility and the diffusion constant at room temperature, the extrapolated barrier at 1.9%, corresponding to the experimental lattice constant at room temperature [42], was used. With an energy barrier of 0.19 eV (the average of the two 2NN barriers) and a hopping distance of the 2NN Fe-Fe distance, the computed electron mobility in bulk BiFeO3 at room temperature is cm2/Vs, the computed diffusion constant is cm2/s. With an estimated uncertainty of 10% in the activation barrier, one obtains to cm2/Vs and to cm2/s.
3.3 Excess electrons at domain walls
3.3.1 Trapping of polarons at domain walls
The atomic structure, ionic ferroelectric polarization profile , total energy change with respect to a polaron position in the domain interior, relative polaron density as a function of temperature (: density in the domain interior), in-plane averaged and smoothened electronic potential , local direction-averaged strain, and octahedral tilt profiles around the three cartesian axes for small electron polarons near these domain walls are shown in Figure 6. The density was calculated with Eq. (1), the ionic polarization with Eq. (5).
The local strain with polaron in the lowest-energy site appears to couple to existing strain variations at the walls, possibly also to tilt variations. The local electronic potential in the pristine system, before polaron formation, exhibits a minimum at or near the lowest-energy polaron site.
The trapping energies for small electron polarons at the domain walls are listed in Table 4.
| DW | 160-atom supercell | 240-atom supercell |
| 71° | -0.063 | -0.051 |
| 109° | -0.075 | -0.065 |
| 180° | -0.044 | -0.054 |
We observe two different trap depths for the two 180° domain walls even though the two domain walls are equivalent by symmetry; this is because the symmetry operation relating first and second wall exchanges the magnetic sublattices. Here only polarons on spin-down sites were considered, the profile for polarons of the other spin was then obtained based on symmetry. The computed trapping energies (-51 meV for the 71°, -65 meV for the 109°, and -54 meV for the 180° domain wall) are similar to each other. The 180° domain wall differs from the 71° and 109° walls in that it attracts small polarons in a plane one layer apart from the layer closest to the wall, and, interestingly, each of the two equivalent 180° walls selectively traps essentially only polarons with a specific spin. The trapping energies were linearly corrected an for underestimated ferroelectric energy gain with PBEsol as the reference, see Table 1 and section 2.2. The short-range potential variations at the doman walls are superimposed on a large-scale potential landscape that depends on the boundary conditions, for example, open- or closed-circuit conditions, surface termination, etc. To estimate the density of excess electrons at the walls, one should consider both short-range and long-range potential. In a previous work [20], the long-range potential at 71° and 109° domain walls was found to consist in a potential step at the wall and a potential slope in the domains that depends on the boundary conditions. For zero potential slope (zero electric field in the domains), only the short-range potential acts on the polarons, and their density should be determined by Eq. 1, as depicted in Figure 6. For closed-circuit conditions, the potential step at the domain walls should create an electric field in the domains. This case is considered in Figure 7, assuming a domain-wall spacing of 150 nm [43, 44] and a polaron density in the dilute limit ( and very small).
3.3.2 Polaron hopping at domain walls
For the 71° domain wall, polaron hopping was investigated along and across the domain wall. When starting from a delocalized-electron configuration at the 71° domain wall, a small electron polaron forms spontaneously in the plane labeled “1” (on the right-hand side of the wall, if the net polarization points to the right, see Figure 2(b)), which is a metastable site. The most stable site is inside the domain wall, in the plane labeled “0”.
Figure 8 shows the energy barriers for electron hops at the 71° domain wall. Only some representative low barriers in each direction are shown. The computed 2NN energy barriers are listed in Table 5. All barriers are depicted and tabulated in the Supporting Information. The hops are projected on the symmetry-adapted coordinates [ wall, (a)], [within the wall, (b)], and [within the wall, , (c)]. For hops between the two trap states in the planes “0” and “1”, the barrier is about 0.2 eV; for hops out of the wall (from plane “0” or “1” to plane “-1”,“2”, or “-2”) the barrier is about 10 to 20% larger. The barriers are invariant with respect to translation by a lattice vector parallel to the wall. Like in the bulk, the barriers slightly increase under tensile strain. The lowest barriers for NN and 2NN hops are similar in size.
| double 2NN, | 2NN, | 2NN, escape | (cm2/Vs) | ||||
| or | from DW | (cm2/s) | |||||
| % | 0.19 | 0.19 | 0.22 | – | – | ||
| 0 | 0.20 | 0.19 | 0.23 | – | – | ||
| 0.34% | 0.20 | 0.20 | 0.23 | – | – | ||
| 1.9% (extrapol.) | 0.21 | 0.21 | 0.25 | ||||
The computed two-dimensional electron mobility inside the domain-wall plane is cm2/Vs, the computed two-dimensional diffusion constant inside the domain wall is cm2/s. Assuming 10% uncertainty in the barriers, one obtains 0.7 to 4 cm2/Vs and 0.2 to 1 cm2/s. The temperature dependence of the computed mobility is depicted in Figure 9. The mobility is dominated by the exponential decay originating in the hopping barrier.
4 Discussion
All three ferroelectric domain walls, the 71°, the 109°, and the 180° walls, trap small polarons on a short range, with a trap depth of about 50 to 65 meV. Depending on the boundary conditions, such as applied voltage or surface termination, a long-range potential exists in the case of the 71° and 109° domain wall, which originates from a potential step at the domain walls and can create an electric field in the domains, whose strength depends on the boundary conditions: the field should be maximal for closed-circuit conditions and zero for open-circuit conditions. For symmetric 180° domain wall, there should be no electric field. The electric field, if present, dominates the potential energy landscape for small polarons if the potential step at the wall is large compared to the trapping energy, which is the case for the 71° domain wall (potential step: 0.13 eV, trap depth: 51 meV). Assuming a resolution of 10 nm for spatially resolved conductivity measurements at domain walls, the increase in polaron density at the domain wall due to the short-range trap should be hidden behind the large density increase due to the long-range trap originating in the potential step at the wall. In the case of open-circuit conditions, the polaron density and conductivity increase at the wall should originate from the short-ranged trapping profile. In this case, only a moderate increase in polaron density and conductivity of about one order of magnitude is to be expected if measured on a ten-nanometer scale, and only at low temperatures, e.g. at 100 K. The electrical conductivity in the domain interior and at the domain walls will be further influenced by atomic point defects, which are not considered here.
Hall measurements for hematite resulted in electron mobilities of about 0.01 to 1 cm2/Vs [45, 46, 47], which is one to two orders of magnitude higher than the here computed mobility in bulk BiFeO3 of to cm2/Vs. Considering that the Fe-Fe distance in BiFeO3 is larger than in hematite, and that NN hops are spin-forbidden in BiFeO3, a lower mobility is to be expected in BiFeO3. In order to obtain the same mobility for BiFeO3 as in hematite, the energy barriers would need to be about 40% smaller, near 0.1 eV instead of 0.2 eV. At elevated temperatures, a combination of electron tunneling through the top of the barrier, smearing out of the traps due to thermal motion, incomplete self-trapping, and occupation of excited self-trapped states should effectively reduce the barriers. At low temperature, the effective barriers should approach those computed here. The here calculated hopping barriers ( eV) are similar to those obtained for PbTiO3 [31], so that the activation energy for the intrinsic electron mobility in BiFeO3 should be comparable to that in PbTiO3.
Experimentally determined activiation energies of electron conduction in BiFeO3 are typically in a similar range or larger (between about 0.1 and 1.3 eV [48, 49, 50, 51, 52]) than the here computed barrier of about 0.2 eV. The lower measured barriers could hence originate from electron polaron hopping, while it seems likely that the experimentally determined higher barriers of several hundred meV originate from defects other than pristine neutral bulk or ferroelectric domain walls, for example from point defects, such as oxygen or bismuth vacancies, or from interface effects, such as Schottky barriers.
In Refs. [13, 53] an activation barrier of eV was found for -type conduction, which is about the energy difference between the electron polaron level and the conduction band minimum computed here. In principle, this activation barrier of 0.7 eV could originate from emission of electrons from the polaron state to the conduction band. This would mean that transport by emission from polarons into delocalized conduction states and subsequent ballistic transport would be more efficient than polaron hopping, inspite of the much higher energy barrier that must initially be overcome. However, there are reasons to believe that the lifetime of thermalized electrons in the conduction band at room temperature should be small: The strong Fe- contribution to the states at the conduction band minimum and the small size of the electron polaron indicate that electron-phonon coupling should be large, which should lead to strong electron-phonon scattering. First-principles molecular dynamics simulations yield a polaron formation time of less than a picosecond, see Supporting Information, similar to hematite [54]. For hematite, transient extreme-ultraviolet spectroscopy yielded “signatures characteristic of electron localization by small polarons […] within 100 fs following photoexcitation” [55]. Hence, it seems more likely that the measured barrier of 0.7 eV belongs to a defect or interface.
Once a small electron polaron has formed, it is unlikely to transform into a delocalized conduction-band electron at room temperature and below since the energy distance between the polaron level and the conduction-band minimum of eV is much larger than the polaron hopping barrier of eV. It seems most likely that small electron hopping should be the dominant intrinsic -type transport process at room temperature and below.
Here it is assumed that excess electrons have been brought into BiFeO3 in some way. For example, charge carriers may originate from point defects, such as oxygen vacancies or aliovalent dopants or impurities, from photoexcitation resulting in bound or dissociated electron-hole pairs, or from an electrical contact.
5 Summary and conclusion
Summarizing, the domain wall planes act as traps for small electron polarons; trapping energies of -51 meV (71° wall), -75 meV (109° wall), and -54 meV (180° wall) were obtained here. The origin of the trapping is likely a coupling of the polaron deformation to existing strain and possibly tilt variations at the domain walls. These trap depths are about two times at room temperature and hence near the boundary between deep and shallow levels. Assuming as the boundary between shallow and deep levels [56], the domain walls act as shallow traps at and above room temperature and become deep traps for temperatures below to K. The atomic and magnetic structure of the 180° domain walls couple in such a way that the walls selectively trap polarons depending on their spin. Pristine domain walls, especially 71° domain walls, can accumulate polarons in a region of tens of nanometers up to densities that are several orders of magnitude larger than in the bulk, depending on the boundary conditions. This is due to a potential step in the wall that can create a long-ranged electric field. For the 109° domain wall, a maximum polaron density of less than one order of magnitude more than in the bulk is possible based on the computed potential step, at symmetric 180° walls there should be no such long-range electric field. The experimentally measured increase in electrical conductivity at the domain walls should essentially originate in a local increase in polaron density at the walls.
For the first time, energy barriers for small electron polaron hopping in BiFeO3 were determined using density-functional theory. Also for the first time, electron polaron hopping at a ferroelectric domain wall was modeled from first principles. Hopping in perfect, pristine BiFeO3 and at pristine, neutral ferroelectric 71° domain walls was considered.
In both bulk and domain-wall plane, the calculated energy barriers for hopping are about 0.2 eV
and slightly increase under tensile strain.
The calculated hopping barrier for electrons in BiFeO3 is similar, and the resulting activation barrier for the electron mobility should be similar
to, e.g., PbTiO3. Using non-adiabatic Marcus theory, an electron mobility of
to cm2/Vs and a computed diffusion constant of 0.2 to 3 cm2/s
are obtained at room temperature both in the bulk and in the domain-wall planes.
Concluding, (a) experimentally found activation barriers of several hundred meV for electrical conductivity in BiFeO3
should have an origin other than electron polaron hopping in pristine BiFeO3 or electron polaron escape from pristine neutral ferroelectric 71° domain walls.
To identify the origin of these large energy barriers is here deferred to future work.
This will be a considerable task, but it should be undertaken to fully understand the limits of electronic conductivity in BiFeO3 and
other transition-metal oxides.
(b) The effect of pristine, neutral ferroelectric domain walls on the -type conductivity at room temperature consists in accumulating excess electrons
and confining their motion to the domain-wall plane.
In the Drude picture (conductivitycarrier densitymobility), the electron density at the wall is enhanced compared to the bulk,
the mobility perpendicular to the wall is slightly reduced compared to the bulk.
This is in line with experimental observations that conductivity is enhanced at ferroelectric domain walls
and in directions parallel to the domain walls.
At temperatures up to room temperature, pristine ferroelectric 71° domain walls can in principle
exhibit a carrier density and an electrical conductivity up to several orders of magnitude larger than in the bulk,
depending on the boundary conditions, due to a potential step of about 0.13 eV in the wall.
For the 109° domain wall, assuming a spatial resolution of ten nanometers,
a carrier density and conductivity enhancement by up to one order of magnitude is possible,
due to a potential step of about 17 meV.
For the symmetric 180° domain wall, the carrier density is enhanced by far less than one order of magnitude,
when averaged on a ten-nm scale.
Accumulation of point defects other than electron polarons, not considered here,
should further modify the potential and carrier density profile of the ferroelectric domain walls.
Supporting Information
Supporting Information is available from the Wiley Online Library or from the author.
Acknowledgements
S. K. thanks Michele Reticcioli for recommending the OMC method. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No. 746964, from a scholarship programme at the University of Jena within the Thuringian Scheme for Young Female Researchers and Young Female Artists funded by the Free State of Thuringia, and from the Volkswagen Stiftung (Momentum) through the project “dandelion”. Computational resources and support were supplied by the HPC cluster ARA of the University of Jena, Germany. Figures were made using Vesta [57] and gnuplot.